If a car is sold for Rs 50000 the profit is 17 percent;what would be profit percentage if sold for Rs 47000?
Answer by jorel555(1290)

If the profit on a 50000 Rs car is 17%, then the price is 1.17 times the cost, or:
1.17cost=Price
cost = 50000/1.17=42735.04
If the selling price is 47000, then the profit is 47000/42735.04=1.0998 cost, or 9.98% profit!!!!!!!!!!!!!!